OpenStudy (anonymous):

how do i work out the problem a-5/10=2a-3/5

OpenStudy (anonymous):

\[a - {5 \over 10 } = 2a - {3 \over 5}\] \[a - { 5 \over 10 } = 2a - { 6 \over 10}\] \[a - 2a = -{6 \over 10 } + { 5 \over 10 }\] \[-a = -{1 \over 10}\] \[a = { 1 \over 10}\]

OpenStudy (radar):

a-5/10 = 2a -3/5 Subtract a from both sides getting -5/10= a -3/5 now add 3/5 to both sides getting -5/10 + 3/5 = a now combine -5/10 +6/10 =a converting 3/5 to 6/10 by multiplying it by 2/2 (1) a=1/10